# A Cosmic Miracle: A Remarkably Luminous Galaxy at Confirmed with JWST

### v1

@@ 4.1 The spectroscopic UV luminosity function at
The bright end of the UV luminosity function (UV LF) at has been a key benchmark for early Universe models in the JWST era. The chasm between theory and observations suggested by GN-z11 (Oesch et al., 2016) at was immediately borne out by photometric samples (e.g., Naidu et al., 2022c; Castellano et al., 2022; Donnan et al., 2023; Harikane et al., 2023b; Finkelstein et al., 2023; Adams et al., 2024) and subsequently confirmed with spectroscopy (e.g., Arrabal Haro et al., 2023; Castellano et al., 2024; Napolitano et al., 2025). The discovery of JADES-GS-z14-0 in a arcmin2 field hinted that bright sources remained abundant even at higher redshifts, where the tension with models may be even more pronounced (e.g., Robertson et al., 2024a; Whitler et al., 2025; Pérez-Gonzålez et al., 2025b; Castellano et al., 2025b).
Here we confirm this picture. The number density implied by MoM-z14 and JADES-GS-z14-0 is shown in Fig. 10. To estimate this LF point we assume MoM-z14 and JADES-GS-z14-0 are the only sources over the UDS, COSMOS, and GOODS-S fields at in the area imaged by the PRIMER and JOF surveys (Donnan et al., 2024; Eisenstein et al., 2023b). Our derived number density â â is in excellent agreement with photometric luminosity functions recently reported at (e.g., Finkelstein et al., 2024; Donnan et al., 2024; Robertson et al., 2024a; Whitler et al., 2025).
@@ 4.2 What makes luminous galaxies shine so bright?
(i) A higher UV variability (), which represents the scatter in UV brightness at fixed mass. The key form of explored in the literature has been âburstyâ star-formation, where galaxies oscillate between periods of quiescence and starbursts on short Myr timescales (e.g., Faucher-GiguĂšre, 2018; Mason et al., 2023; Sun et al., 2023; Nikopoulos & Dayal, 2024; Endsley et al., 2024; Looser et al., 2024; Kravtsov24). While indeed, galaxies like MoM-z14 show a rising SFH, constraining the magnitude of this rise purely from rest-UV SEDs is challenging (see the posteriors in Fig. 6; c.f. Harikane et al. 2025; Kokorev et al. 2025b). We also note that the required to match the LF is much higher, mag (e.g., Shen et al., 2023, 2024), compared to the mag inferred out to from galaxy clustering (e.g., Shuntov et al., 2025).
@@
(ii) A higher star-formation efficiency (SFE), representing the conversion rate between gas and stars. For example, Dekel et al. (2023, 2025); Li et al. (2024c) envision feedback-free star-formation wherein dense gas () at low metallicity () may collapse into stars at a rate much faster than can be regulated by supernovae or stellar winds (see also Boylan-Kolchin 2025). These ISM conditions are not implausible in light of âblack hole starsâ with (e.g., Naidu et al., 2025; de Graaff et al., 2025) thought to power Little Red Dots â LRDs occur at a similar rate as the luminous galaxies (; e.g., Matthee et al. 2024; Akins et al. 2024; Kokorev et al. 2024). Indeed, MoM-z14 does show hints of the required ISM conditions for this scenario (Fig. 7) â this can be further tested with high-resolution spectroscopy to resolve density-sensitive multiplets (e.g., Nâiii]; Maiolino et al. 2024).
@@
(iii) Modifications to CDM that enhance the abundance and accretion rate of early halos (e.g., Liu & Bromm, 2022; Huang et al., 2023; Gouttenoire et al., 2024). For instance, in Early Dark Energy (EDE), an increased early expansion relative to CDM decreases the physical sound horizon measured in CMB, resulting in shifts of the inferred cosmological parameters. This solves the Hubble tension with a side effect of enhancing dark matter halo abundance at high redshifts (see Poulin et al. 2023 for a recent review). Clustering is a promising test of whether luminous galaxies are resident in truly massive halos (e.g., Muñoz et al., 2023; Gelli et al., 2024; Eilers et al., 2024) â in the EDE scenario one would expect a higher halo mass on average relative to say CDM with a high . At the moment, we do not find any obviously associated sources around MoM-z14, but much deeper imaging and spectroscopy will be required to rule out any companions. No spectroscopically associated sources or overdensities have been reported for other luminous sources either (though see Tacchella et al. 2023a for possible photometric companions around GNz11).
@@
(iv) Boosting the light to mass ratio relative to lower-redshift galaxies. Examples include decreasing dust attenuation with redshift (e.g., Ferrara et al., 2023), a top-heavy IMF (e.g., Yung et al., 2024; Hutter et al., 2025; Lu et al., 2025), or a high incidence of AGN (e.g., Maiolino et al., 2024; CalabrĂČ et al., 2024; Natarajan et al., 2024). MoM-z14 does show hints of a blue slope consistent with little dust, and a chemical abundance pattern consistent with very massive stars that may emerge under a top-heavy IMF (see §4.3). The presence of strong UV lines in the prism spectrum guarantee that higher resolution spectroscopy will be fruitful in drastically shrinking the posteriors in Fig. 7 (e.g., of and [N/C]) to test these scenarios.
@@ 4.3 Insights from the Milky Way into MoM-z14, early bright galaxies, and supermassive black holes
Insight into the underlying physics of strong nitrogen emitters may be found in the Milky Way, where nitrogen enhancement is a characteristic feature of globular cluster (GC) stars â see Bastian & Lardo (2018) and Gratton et al. (2019) for recent reviews. Furthermore, super-solar nitrogen is also observed in 2â4% of the most metal-poor stars that formed in the Milky Way prior to the formation of the disk(s) at (the âAuroraâ or âproto-galaxyâ component of the Milky Way; Conroy et al. 2022; Rix et al. 2022; Belokurov & Kravtsov 2022; Chandra et al. 2024; Semenov et al. 2024). Note that this is a distinct component of the stellar halo, the majority of which is comprised of debris from accreted dwarf galaxies (e.g., Belokurov et al., 2018; Helmi et al., 2018; Naidu et al., 2020a). By using nitrogen as a proxy for formation in dense clusters, of star-formation at in the Milky Way (i.e., in Aurora) is estimated to have occurred within such bound structures (e.g., Horta et al., 2021; Belokurov & Kravtsov, 2023). Intriguingly, the fraction of N-enhanced stars then rapidly falls off towards higher metallicities () by few orders of magnitude as the Milky Wayâs disk begins spinning up (e.g., Belokurov & Kravtsov, 2023, 2024; Kane et al., 2025). Indeed, direct observations of lensed galaxies at are direct reflections of this archeological inference (e.g., Fujimoto et al., 2024; Adamo et al., 2024). Therefore, the N-enhancement in MoM-z14 and other high- sources seems far from peculiar, and perhaps reflects the generic mode of star-formation in very dense clusters prevalent at these epochs (e.g., Ma et al., 2020).
@@
As for the underlying physical picture, these primordial dense clusters may be hotbeds for peculiar stellar populations and unique ISM conditions. For example, the rate of rapid runaway collisions in massive GCs is expected to be significant, and may help produce âvery massive starsâ (VMS, ; e.g., Vink 2024) or âsupermassive starsâ (SMS, ; e.g., Denissenkov & Hartwick 2014; Gieles et al. 2018; Vergara et al. 2025). These stars would then leave behind an enriched ISM with characteristic chemical abundance patterns (from e.g., hot H burning in the CNO cycle; Charbonnel et al. 2023).
@@
Further, beyond abundances, with objects like SMS and VMS, the top-heavy IMF and drastically altered mass-to-light ratio (e.g., Martins et al., 2020) would help reconcile theoretical and observed UV LFs. The we report for MoM-z14 then may be an upper limit on the stellar mass, implying massive () GCs may account for all the light. This may also be linked to the compactness of N-emitters â if only a few GCs dominate the light, they may produce a light profile with a very small effective radius (Fig. 5). The stellar surface density we infer in MoM-z14 ( ) is comparable to Milky Way GCs, even if the stellar mass is over-estimated by 1-2 dex (e.g., Hopkins et al., 2010). The hard ionizing spectrum (§3.2.3) could also be naturally explained by an IMF that is dominated by these highly ionizing massive stars (e.g. Martins et al., 2025). What then of the extended sources such as JADES-GS-z14-0 comprising the upper branch in Fig. 5? These sources may represent a mode of star-formation corresponding to the of nitrogen-poor stars forming in the Milky Way in these early epochs.
@@
As for the connection between UV-bright galaxies, N-rich LRDs (e.g., Labbe et al., 2024), and broad-line AGN (e.g., Isobe et al., 2025), perhaps the remains of VMS and SMS mark the beginnings of gas-enshrouded supermassive black holes (âblack hole starsâ) that form deeply embedded in this dense environment (e.g., Ji et al., 2025; Naidu et al., 2025; de Graaff et al., 2025; Rusakov et al., 2025; Taylor et al., 2025; Marques-Chaves et al., 2024b). It is tempting to speculate that the drop in LRD number-density by an order of magnitude below (e.g., Ma et al., 2025; Euclid Collaboration et al., 2025) is linked to the similar drop in the occurrence of N-enriched stars in the Milky Way at these redshifts. These promising, yet preliminary connections between the highest redshift galaxies and the local archaeological record hinge on detailed chemical abundances, motivating investments in deep, high-resolution spectroscopy.
@@ 5 Summary & Outlook
This paper presented first results from the âMirage or Miracleâ (MoM) survey that we designed to spectroscopically test the abundance and nature of luminous galaxy candidates at . Here we introduced MoM-z14, a remarkably bright () source at whose discovery pushes the cosmic frontier to million years after the Big Bang. We find the following:
âą
The NIRCam imaging of MoM-z14 shows a robust dropout signature (), which the follow-up NIRSpec prism spectrum reveals is due to a sharp Ly break at . Remarkably for a galaxy at this early epoch, the prism spectrum displays multiple rest-UV emission lines (Câiv, Câiii], Nâiv, Nâiii, Heâii+Oâiii]). This allows for a rather precise redshift determination () while also providing a unique opportunity for detailed physical characterization. [Figs. 1, 2, Table 2.2, §3.1]
- âą
- âą
âą
@@
âą
MoM-z14 is a strong Nâiv]Ă Â emitter, adding the highest redshift example yet to this emerging class of sources that now includes a collection of luminous Little Red Dots, broad-line AGN, and extremely compact star-forming galaxies. In fact, pending confirmation with high-resolution spectroscopy, MoM-z14 may rank among the most nitrogen-enhanced sources discovered with JWST yet ([N/C]). It adds further evidence for a size-chemistry bimodality at , wherein extended sources tend to be nitrogen weak while compact sources are strong N emitters. [Figs. 5, 8, §3.2.3]
âą
We interpret MoM-z14 and N-emitters through Galactic archaeology, connecting their abundance patterns to the most ancient stars born in the Milky Way at (âAuroraâ/the âproto-galaxyâ) as well as to globular clusters. The N-enhancement, brightness, hard ionizing spectra, stellar density, morphology, redshift dependence, and black hole fraction of these sources may be linked to GC-like environments wherein runaway collisions may produce extraordinary objects such as supermassive stars. [Figs. 5, 8, §4.3]
That the number density of luminous galaxies evolves only gradually between and is now on firm spectroscopic footing. The good fortune of inhabiting a Universe teeming with GN-z11s means remarkably luminous galaxies in the hundreds may be within the grasp of the Roman Space Telescope. JWST itself appears poised to drive a series of great expansions of the cosmic frontier â previously unimaginable redshifts, approaching the era of the very first stars, no longer seem far away.
@@ Acknowledgments
âMirage or Miracleâ is but the latest link in a long chain of surveys that have built COSMOS into a premier extragalactic legacy field. We are thankful to all the teams who have contributed to this legacy, particularly those mentioned in §3 for leading recent JWST programs whose imaging we have incorporated in our analysis. We are grateful to Vasily Belokurov for help in compiling the Milky Way reference sample featured in Fig 8. We thank Danielle Berg for sharing a highly complete, highly decimalized NUV vacuum line list. We are grateful to our programâs NIRSpec reviewer, Dan Coe, and program coordinator, Allison Vick, for valuable input on our MSA design. We acknowledge illuminating conversations with Risa Wechsler and Chao-Lin Kuo about early reionization. RPN thanks Neil Pappalardo and Jane Pappalardo for their generous support of the MIT Pappalardo Fellowships in Physics, and for their enthusiasm and encouragement for seeking galaxies at the highest redshifts.
@@
RPN acknowledges funding from JWST program GO-5224. Support for this work was provided by NASA through the NASA Hubble Fellowship grant HST-HF2-51515.001-A awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555. This work has received funding from the Swiss State Secretariat for Education, Research and Innovation (SERI) under contract number MB22.00072, as well as from the Swiss National Science Foundation (SNSF) through project grant 200020_207349. Funded by the European Union (ERC, AGENTS, 101076224 and HEAVYMETAL, 101071865). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. The Cosmic Dawn Center (DAWN) is funded by the Danish National Research Foundation under grant DNRF140. This work has also been supported by JSPS KAKENHI Grant Number 23H00131. HA acknowledges support from CNES, focused on the JWST mission, and the Programme National Cosmology and Galaxies (PNCG) of CNRS/INSU with INP and IN2P3, co-funded by CEA and CNES. SB is supported by the UK Research and Innovation (UKRI) Future Leaders Fellowship [grant number MR/V023381/1]. R.D. acknowledges support from the INAF GO 2022 grant âThe birth of the giants: JWST sheds light on the build-up of quasars at cosmic dawnâ and by the PRIN MUR â2022935STWâ, RFF M4.C2.1.1, CUP J53D23001570006 and C53D23000950006.
@@
Computations supporting this paper were run on MITâs Engaging cluster. This publication made use of the NASA Astrophysical Data System for bibliographic information. Some of the data products presented herein were retrieved from the Dawn JWST Archive (DJA). DJA is an initiative of the Cosmic Dawn Center (DAWN), which is funded by the Danish National Research Foundation under grant DNRF140. Software used in developing this work includes: matplotlib (Hunter, 2007), jupyter (Kluyver et al., 2016), IPython (PĂ©rez & Granger, 2007), numpy (Oliphant, 2015), scipy (Virtanen et al., 2020), TOPCAT (Taylor, 2005), and Astropy (Astropy Collaboration et al., 2013).

### v2

@@ 4.1 The spectroscopic UV luminosity function at
The bright end of the UV luminosity function (UV LF) at has been a key benchmark for early Universe models in the JWST era. The chasm between theory and observations suggested by GN-z11 (Oesch et al., 2016) at was immediately borne out by photometric samples (e.g., Naidu et al., 2022c; Castellano et al., 2022; Donnan et al., 2023; Harikane et al., 2023b; Finkelstein et al., 2023; Adams et al., 2024) and subsequently confirmed with spectroscopy (e.g., Arrabal Haro et al., 2023; Castellano et al., 2024; Napolitano et al., 2025). The discovery of JADES-GS-z14-0 in a arcmin2 field hinted that bright sources remained abundant even at higher redshifts, where the tension with models may be even more pronounced (e.g., Robertson et al., 2024a; Whitler et al., 2025; Pérez-González et al., 2025b; Castellano et al., 2025b).
Here we confirm this picture. The number density implied by MoM-z14 and JADES-GS-z14-0 is shown in Fig. 11. To estimate this LF point we assume MoM-z14 and JADES-GS-z14-0 are the only sources over the UDS, COSMOS, and GOODS-S fields at in the area imaged by the PRIMER and JOF surveys (Donnan et al., 2024; Eisenstein et al., 2023a). The survey area used for this calculation is 136 (PRIMER-COSMOS; Donnan et al. 2024) + 243 (PRIMER-UDS; Donnan et al. 2024) + 9 arcmin2 (JOF; Robertson et al. 2024a). The differential co-moving volume at for this survey area is 462,648 Mpc3.. Our derived number density – – is in excellent agreement with photometric luminosity functions recently reported at (e.g., Finkelstein et al., 2024; Donnan et al., 2024; Robertson et al., 2024a; Whitler et al., 2025).
@@ 4.2 What makes luminous galaxies shine so bright?
(i) A higher UV variability (), which represents the scatter in UV brightness at fixed mass. The key form of explored in the literature has been “bursty” star-formation, where galaxies oscillate between periods of quiescence and starbursts on short Myr timescales (e.g., Faucher-Giguère, 2018; Mason et al., 2023; Sun et al., 2023; Nikopoulos and Dayal, 2024; Endsley et al., 2024; Looser et al., 2024; Kravtsov and Belokurov, 2024). While indeed, galaxies like MoM-z14 show a rising SFH, constraining the magnitude of this rise purely from rest-UV SEDs is challenging (see the posteriors in Fig. 6; c.f. Harikane et al. 2025; Kokorev et al. 2025b). We also note that the required to match the LF is much higher, mag (e.g., Shen et al., 2023, 2024), compared to the mag inferred out to from galaxy clustering (e.g., Shuntov et al., 2025).
@@
(ii) A higher star-formation efficiency (SFE), representing the conversion rate between gas and stars. For example, Dekel et al. (2023, 2025); Li et al. (2024c) envision feedback-free star-formation wherein dense gas () at low metallicity () may collapse into stars at a rate much faster than can be regulated by supernovae or stellar winds (see also Boylan-Kolchin 2025). These ISM conditions are not implausible in light of “black hole stars” with (e.g., Naidu et al., 2025; de Graaff et al., 2025) thought to power Little Red Dots – LRDs occur at a similar rate as the luminous galaxies (; e.g., Matthee et al. 2024; Akins et al. 2024; Kokorev et al. 2024). Indeed, MoM-z14 does show hints of the required ISM conditions for this scenario (Fig. 7) – this can be further tested with high-resolution spectroscopy to resolve density-sensitive multiplets (e.g., N iii]; Maiolino et al. 2024).
@@
(iii) Modifications to CDM that enhance the abundance and accretion rate of early halos (e.g., Liu and Bromm, 2022; Huang et al., 2023; Parashari and Laha, 2023; Gouttenoire et al., 2024). For instance, in Early Dark Energy (EDE), an increased early expansion relative to CDM decreases the physical sound horizon measured in CMB, resulting in shifts of the inferred cosmological parameters. This solves the Hubble tension with a side effect of enhancing dark matter halo abundance at high redshifts (see Poulin et al. 2023 for a recent review). Clustering is a promising test of whether luminous galaxies are resident in truly massive halos (e.g., Muñoz et al., 2023; Gelli et al., 2024; Eilers et al., 2024) – in the EDE scenario one would expect a higher halo mass on average relative to say CDM with a high . At the moment, we do not find any obviously associated sources around MoM-z14, but much deeper imaging and spectroscopy will be required to rule out any companions. No spectroscopically associated sources or overdensities have been reported for other luminous sources either (though see Tacchella et al. 2023a for possible photometric companions around GNz11).
@@
(iv) Boosting the light to mass ratio relative to lower-redshift galaxies. Examples include decreasing dust attenuation with redshift (e.g., Ferrara et al., 2023), a top-heavy IMF (e.g., Yung et al., 2024; Hutter et al., 2025; Lu et al., 2025), or a high incidence of AGN (e.g., Maiolino et al., 2024; Calabrò et al., 2024; Natarajan et al., 2024). MoM-z14 does show hints of a blue slope consistent with little dust, and a chemical abundance pattern consistent with very massive stars that may emerge under a top-heavy IMF (see §4.3). The presence of strong UV lines in the prism spectrum guarantee that higher resolution spectroscopy will be fruitful in drastically shrinking the posteriors in Fig. 7 (e.g., of and [N/C]) to test these scenarios.
@@ 4.3 Insights from the Milky Way into MoM-z14, early bright galaxies, and supermassive black holes
Insight into the underlying physics of strong nitrogen emitters may be found in the Milky Way, where nitrogen enhancement is a characteristic feature of globular cluster (GC) stars – see Bastian and Lardo (2018) and Gratton et al. (2019) for recent reviews. Furthermore, super-solar nitrogen is also observed in 2–4% of the most metal-poor stars that formed in the Milky Way prior to the formation of the disk(s) at (the “Aurora” or “proto-galaxy” component of the Milky Way; Conroy et al. 2022; Rix et al. 2022; Belokurov and Kravtsov 2022; Chandra et al. 2024; Semenov et al. 2024). Note that Aurora is a distinct in-situ (born in the Milky Way) component of the stellar halo, which is different from the majority of halo stars that arise from the debris of accreted dwarf galaxies (e.g., Belokurov et al., 2018; Helmi et al., 2018; Naidu et al., 2020a). By using nitrogen as a proxy for formation in dense clusters, of star-formation at in the Milky Way (i.e., in Aurora) is estimated to have occurred within such bound structures (e.g., Horta et al., 2021; Belokurov and Kravtsov, 2023). Intriguingly, the fraction of N-enhanced stars then rapidly falls off towards higher metallicities () by few orders of magnitude as the Milky Way’s disk begins spinning up (e.g., Belokurov and Kravtsov, 2023, 2024; Kane et al., 2025). Indeed, direct observations of lensed galaxies at are direct reflections of this archeological inference (e.g., Fujimoto et al., 2024; Adamo et al., 2024). Therefore, the N-enhancement in MoM-z14 and other high- sources seems far from peculiar, and perhaps reflects the generic mode of star-formation in very dense clusters prevalent at these epochs (e.g., Ma et al., 2020).
@@
As for the underlying physical picture, these primordial dense clusters may be hotbeds for peculiar stellar populations and unique ISM conditions. For example, the rate of rapid runaway collisions in massive GCs is expected to be significant, and may help produce “very massive stars” (VMS, ; e.g., Vink 2024) or “supermassive stars” (SMS, ; e.g., Denissenkov and Hartwick 2014; Gieles et al. 2018; Vergara et al. 2025). These stars would then leave behind an enriched ISM with characteristic chemical abundance patterns (from e.g., hot H burning in the CNO cycle; Charbonnel et al. 2023).
@@
Further, beyond abundances, with objects like SMS and VMS, the top-heavy IMF and drastically altered mass-to-light ratio (e.g., Martins et al., 2020) would help reconcile theoretical and observed UV LFs. The we report for MoM-z14 then may be an upper limit on the stellar mass, implying massive () GCs may account for all the light. This may also be linked to the compactness of N-emitters – if only a few GCs dominate the light, they may produce a light profile with a very small effective radius (Fig. 5). The stellar surface density we infer in MoM-z14 ( ) is comparable to Milky Way GCs, even if the stellar mass is over-estimated by 1-2 dex (e.g., Hopkins et al., 2010). The hard ionizing spectrum (§3.2.3) could also be naturally explained by an IMF that is dominated by these highly ionizing massive stars (e.g. Martins et al., 2025). What then of the extended sources such as JADES-GS-z14-0 comprising the upper branch in Fig. 5? These sources may represent a mode of star-formation corresponding to the of nitrogen-poor stars forming in the Milky Way in these early epochs.
@@
As for the connection between UV-bright galaxies, N-rich LRDs (e.g., Labbe et al., 2024), and broad-line AGN (e.g., Isobe et al., 2025), perhaps the remains of VMS and SMS mark the beginnings of gas-enshrouded supermassive black holes (“black hole stars”) that form deeply embedded in this dense environment (e.g., Ji et al., 2025; Naidu et al., 2025; de Graaff et al., 2025; Rusakov et al., 2025; Taylor et al., 2025). It is tempting to speculate that the drop in LRD number-density by an order of magnitude below (e.g., Ma et al., 2025; Euclid Collaboration et al., 2025) is linked to the similar drop in the occurrence of N-enriched stars in the Milky Way at these redshifts. These promising, yet preliminary connections between the highest redshift galaxies and the local archaeological record hinge on detailed chemical abundances, motivating investments in deep, high-resolution spectroscopy.
@@ 5 Summary & Outlook
This paper presented first results from the “Mirage or Miracle” (MoM) survey that we designed to spectroscopically test the abundance and nature of luminous galaxy candidates at . Here we introduced MoM-z14, a remarkably bright () source at whose discovery pushes the cosmic frontier to million years after the Big Bang. We find the following:
The NIRCam imaging of MoM-z14 shows a robust dropout signature (), which the follow-up NIRSpec prism spectrum reveals is due to a sharp Ly break at . Remarkably for a galaxy at this early epoch, the prism spectrum displays multiple rest-UV emission lines (C iv, C iii], N iv, N iii, He ii+O iii]). This allows for a rather precise redshift determination () while also providing a unique opportunity for detailed physical characterization. [Figs. 1, 2, Table 2.2, §3.1]
@@
MoM-z14 is a strong N iv]Å emitter, adding the highest redshift example yet to this emerging class of sources that now includes a collection of luminous Little Red Dots, broad-line AGN, and extremely compact star-forming galaxies. In fact, pending confirmation with high-resolution spectroscopy, MoM-z14 may rank among the most nitrogen-enhanced sources discovered with JWST yet ([N/C]). It adds further evidence for a size-chemistry bimodality at , wherein extended sources tend to be nitrogen weak while compact sources are strong N emitters. [Figs. 5, 8, §3.2.3]
We interpret MoM-z14 and N-emitters through Galactic archaeology, connecting their abundance patterns to the most ancient stars born in the Milky Way at (“Aurora”/the “proto-galaxy”) as well as to globular clusters. The N-enhancement, brightness, hard ionizing spectra, stellar density, morphology, redshift dependence, and black hole fraction of these sources may be linked to GC-like environments wherein runaway collisions may produce extraordinary objects such as supermassive stars. [Figs. 5, 8, §4.3]
That the number density of luminous galaxies evolves only gradually between and is now on firm spectroscopic footing. The good fortune of inhabiting a Universe teeming with GN-z11s means remarkably luminous galaxies in the hundreds may be within the grasp of the Roman Space Telescope. JWST itself appears poised to drive a series of great expansions of the cosmic frontier – previously unimaginable redshifts, approaching the era of the very first stars, no longer seem far away.
@@ Acknowledgments
We thank the two anonymous referees for their insightful comments that have strengthened this work. “Mirage or Miracle” is but the latest link in a long chain of surveys that have built COSMOS into a premier extragalactic legacy field. We are thankful to all the teams who have contributed to this legacy, particularly those mentioned in §3 for leading recent JWST programs whose imaging we have incorporated in our analysis. We are grateful to Vasily Belokurov for help in compiling the Milky Way reference sample featured in Fig 8. We thank Danielle Berg for sharing a highly complete, highly decimalized NUV vacuum line list. We are grateful to our program’s NIRSpec reviewer, Dan Coe, and program coordinator, Allison Vick, for valuable input on our MSA design. We acknowledge illuminating conversations with Risa Wechsler and Chao-Lin Kuo about early reionization. RPN thanks Neil Pappalardo and Jane Pappalardo for their generous support of the MIT Pappalardo Fellowships in Physics, and for their enthusiasm and encouragement for seeking galaxies at the highest redshifts.
@@
RPN acknowledges funding from JWST program GO-5224. Support for this work was provided by NASA through the NASA Hubble Fellowship grant HST-HF2-51515.001-A awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555. This work has received funding from the Swiss State Secretariat for Education, Research and Innovation (SERI) under contract number MB22.00072, as well as from the Swiss National Science Foundation (SNSF) through project grant 200020_207349. Funded by the European Union (ERC, AGENTS, 101076224 and HEAVYMETAL, 101071865). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. The Cosmic Dawn Center (DAWN) is funded by the Danish National Research Foundation under grant DNRF140. This work has also been supported by JSPS KAKENHI Grant Number 23H00131. HA acknowledges support from CNES, focused on the JWST mission, and the Programme National Cosmology and Galaxies (PNCG) of CNRS/INSU with INP and IN2P3, co-funded by CEA and CNES. HA is supported by the French National Research Agency (ANR) under the project FIRSTGAL, grant number ANR-24-CE31-0838. SB is supported by the UK Research and Innovation (UKRI) Future Leaders Fellowship [grant number MR/V023381/1]. R.D. acknowledges support from the INAF GO 2022 grant “The birth of the giants: JWST sheds light on the build-up of quasars at cosmic dawn” and by the PRIN MUR “2022935STW”, RFF M4.C2.1.1, CUP J53D23001570006 and C53D23000950006.
@@
Computations supporting this paper were run on MIT’s Engaging cluster. This publication made use of the NASA Astrophysical Data System for bibliographic information. Some of the data products presented herein were retrieved from the Dawn JWST Archive (DJA). DJA is an initiative of the Cosmic Dawn Center (DAWN), which is funded by the Danish National Research Foundation under grant DNRF140. Software used in developing this work includes: matplotlib (Hunter, 2007), jupyter (Kluyver et al., 2016), IPython (Pérez and Granger, 2007), numpy (Oliphant, 2015), scipy (Virtanen et al., 2020), TOPCAT (Taylor, 2005), and Astropy (Astropy Collaboration et al., 2013).